OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. y - 2 = -1/4 (x + 3); (-3, 5) a. y = -3x -5 b. y = 1/4x - 6 c. y = 4x + 17 d. y = 4x + 5

OpenStudy (anonymous):

Hi there! To begin, can you put the equation you were given into y = mx + b form?

OpenStudy (anonymous):

can you just tell me the answer please? I don't have time right now I need a quick answer

OpenStudy (anonymous):

C

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

You may go, but I'll explain it to you anyways, using a quick and easy elimination method, which is a good test habit. For a line to be perpendicular to another, its slope needs to be the negative inverse of the other line. Your given slope is -1/4, the inverse if -1/4 is -4, the negative of -4 is 4. Therefore your slope is 4. The only two equations with 4 as the slope are C and D. We want the line to go through (-3, 5). Let's use substitution and see which equation passes through this point. I am using C: (5) = 4(-3) + 17 (5) = -12 + 17 5 = 5 C is the answer.